OpenStudy (anonymous):

The 1906 San Francisco earthquake had a magnitude of 7.9 on the MMS scale. At the same time there was an earthquake with magnitude 5 that caused only minor damage. How many times more intense was the San Francisco earthquake than the South American one?

OpenStudy (anonymous):

What is formula for relationship between MMS scale and intensity? {[Logarithmic]

OpenStudy (anonymous):

log(A/B)

OpenStudy (anonymous):

The way I was trying to solve it is 10^R = Richter scale 10^7.9/10^5 = 10^2.9 = 794.33 but I am wrong

OpenStudy (usukidoll):

I don't like this question. Tokyo just had an earthquake -_- bad taste whoever wrote this

OpenStudy (anonymous):

I think I have part of the answer but I am not sure how to determine how much stronger it is.

OpenStudy (mathstudent55):

\(\LARGE dE = 10^{\frac{3}{2}(m_1 - m_2)} \) dE = difference in energy \(m_1\), \(m_2\) = intensities of earthquakes

OpenStudy (anonymous):

I got that far but I need to figure out how much more intense the first earthquake is from the second earthquake. Is there another formula?

OpenStudy (mathstudent55):

Substitute the variables with numbers and solve for dE. \(m_1 = 7.9\), \(m_2 = 5\)

OpenStudy (anonymous):

I am trying to figure out how to enter it on the calculator.

OpenStudy (anonymous):

TI-83

OpenStudy (mathstudent55):

7.9 - 5 = 2.9 1.5 * 2.9 = 4.35 10^4.35 = 22,387

OpenStudy (anonymous):

I put it in as 10*1.5^(7.9-5)

OpenStudy (anonymous):

It returns 32.4089269

OpenStudy (anonymous):

I am probably off with the ()

OpenStudy (mathstudent55):

This is what it should be: 10^[1.5*(7.9-5)]

OpenStudy (anonymous):

On the calculator, I put (10^(1.5)*(7.9-5))

OpenStudy (anonymous):

I have a difficult time to get it right on the calculator so need all the practice

OpenStudy (mathstudent55):

No.

OpenStudy (mathstudent55):

Put parentheses around the entire exponent part. 10^(1.5*(7.9-5))

OpenStudy (anonymous):

5623.41

OpenStudy (anonymous):

It should be 22,387

OpenStudy (mathstudent55):

That was my answer a few responses ago.

OpenStudy (anonymous):

Yes, I am trying to reproduce your answer on the calculator. Thanks for being patient with me

OpenStudy (mathstudent55):

Do it in two steps. First calculate the exponent. 1.5 * (7.9 - 5) What do you get?

OpenStudy (anonymous):

4.35

OpenStudy (mathstudent55):

Great. Now do the power: 10^4.35

OpenStudy (anonymous):

22387

OpenStudy (mathstudent55):

That's it.

OpenStudy (anonymous):

Ok, so it is not possible to input that in one step on the calculator and come up with the same answer? I often try to put it all in one equation and return the answer

OpenStudy (mathstudent55):

You should be able to.

OpenStudy (mathstudent55):

10^(1.5*(7.9-5)) =

OpenStudy (anonymous):

That was it.

OpenStudy (anonymous):

I put in an extra parenthesis and that gave me a different answer.

OpenStudy (mathstudent55):

ok

OpenStudy (anonymous):

10^(1.5*(7.9-5)

OpenStudy (anonymous):

I left off the second parenthesis which is kind of weird.

OpenStudy (anonymous):

Thank you for taking the time to help.

OpenStudy (mathstudent55):

You're welcome.